From mar@roe.ac.uk Fri Apr  4 17:21:24 2014
Date: Wed, 2 Apr 2014 17:16:34 +0100
From: Mike Read <mar@roe.ac.uk>
To: VDFS Science Archive Team -- Clive Davenhall <acd@roe.ac.uk>,
     Dave Morris <dmr@roe.ac.uk>, Eckhard Sutorius <etws@roe.ac.uk>,
     Mark Holliman <msh@roe.ac.uk>, Nicholas Cross <njc@roe.ac.uk>,
     Nigel Hambly <nch@roe.ac.uk>, Rob Blake <rpb@roe.ac.uk>,
     Ross Collins <rsc@roe.ac.uk>, Stelios Voutsinas <stv@roe.ac.uk>
Cc: CCs for VDFS -- Alastair Edge <Alastair.Edge@durham.ac.uk>,
     Andrew Lawrence <al@roe.ac.uk>, Bob Mann <rgm@roe.ac.uk>,
     Dante Minniti <dante@astro.puc.cl>, James Dunlop <jsd@roe.ac.uk>,
     Jim Emerson <j.p.emerson@qmul.ac.uk>, Keith Noddle <keith@keithnoddle.org>,
     Lorenzo Rimoldini <Lorenzo.Rimoldini@unige.ch>,
     Maria-Rosa Cioni <mcioni@aip.de>, matt.jarvis@astro.ox.ac.uk,
     Mike Irwin <mike@ast.cam.ac.uk>,
     Nigel Metcalfe <nigel.metcalfe@durham.ac.uk>,
     Norman Gray <norman@astro.gla.ac.uk>, Phil Lucas <pwl@star.herts.ac.uk>,
     Richard McMahon at IOA <rgm@ast.cam.ac.uk>,
     Simon Dye <Simon.Dye@nottingham.ac.uk>,
     Stephen Warren <s.j.warren@ic.ac.uk>, Tom Kerr <tkerr@jach.hawaii.edu>,
     Tom Shanks <tom.shanks@durham.ac.uk>


WFAU VDFS Science Archives progress meeting minutes, 28 March 2014

Present: RGM, NJC, MAR, ETWS, RPB
Apologies: NCH, KTN, AL, ACD, RSC, MSH

DoNM: 10am Friday 11 April 2014 in the Villas meeting room,
    ACD in the chair (room is booked).


Actions discharged this time:
----------------------------

ACTION: ACD to discuss with RGM the available budget for hardware
     purchases.
Done, should be OK for next 12 months.

ACTION: ALL: make sure that all forthcoming absences during the
     next couple of months are entered into the 'staff absences' Wiki
     page.
Done, WFAU quarterly set for 28th April.

ACTION: RGM to check with AL on progress of negotiations with the
    new owners of UKIRT.
Done, discussions are ongoing but indications are positive.


Actions partly discharged but continuing:
----------------------------------------

ACTION: MSH to e-mail RGM suggested improvements to C2.
It was thought that ACD would speak to Ewan about C2 but
as neither ACD or MSH were present this should be raised at the next 
meeting.


Actions carried forward from 14/3/2014 meeting:
----------------------------------------------

ACTION: ACD, RPB to formulate a policy for backups and for
     checking that data could be recovered from the backups.
Continues.  Notes on a proposed policy are available on the Twiki
     and discussion is still in progress.
ACTION: RPB to put all prices of additional items associated
     with  the policy on the Twiki pages.
Continues
ACTION: All to read the Twiki pages:
     http://apache.roe.ac.uk/twiki/bin/view/WFAU/BackupReview
Some edits have been made. 
Continues

ACTION: RSC to investigate modifying CU7 to recover ATLAS VST
     unmatched sources around frameset boundaries.
Continues.  Investigations had begun and a strategy had been worked
     out.  However, this action now has lower priority.  Nonetheless
     a fix should be put in place before the ATLAS meeting on 14-16
     April and it should be followed by a consortium release shortly
     after the meeting.

ACTION: ALL to add/maintain entries on
     http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines
Continues.

ACTION: RGM to investigate the size and content (how does it differ
     from DR9?) of DR10 to see if it's worth us hosting and whether
     it is necessary to keep all of the other releases.
Continues.  RGM had started a wiki page for WFAU SDSS strategy
     (http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan).
     However, it needs more work and in particular RGM still had to
     determine which WFAU archives had cross-matches with which SDSS
     releases.

ACTION: NJC to sort out the magnitude limits for V1.3 data.
Continues.  The software was largely written, but NJC had not had
     the time to finish and test it.

ACTION: ETWS, with RPB, to develop an automated helper script to
     keep release DB files links updated.
Continues: this one still has low priority.


Specific points and new actions:
-------------------------------

Project management:

Two new posts, one a replacement for Alex, are to be recruited to work 
on GAIA.


WFCAM, VISTA and VST updates:

Apparently the UKIRT handover ceremony did not take place.


Comments and issues arising from CASU minutes:

Nothing to report


Networking:

No problems.


WSA/VSA/OSA/GES Operations:

ETWS reported that in terms of CU4
   - VISTA: (non VVV): complete to Jan 14
   - WFCAM: complete to  Dec 13
   - VST: complete to Jan 14

VVV has been CU4'ed up to Sept 13 (P91).

The latest VHS release had finished outgesting but RPB noted that 
space still needed to be freed up on a public server to host it.

VMC - ppErrBits and tilePawprints had been run.

VVV - tempDB had run out of space crashing objID updates and breaking
replication.

GES - the iDR2 data have been transferred and then re-transferred.


Hardware and Systems:

The new NAS box and processing node had arrived. Eclipse have also
loaned us some 10Gb switches and cards for testing.

The new test system (Nytro RAID card) is installed and being used.


Survey Data Release:

VMC and VVV release files had been uploaded to ESO.

It was noted that the GPS catalogue tables in DR8 could now be made public.
ACTION: RPB to make GPS public in DR8.


Software:

NJC reported working on improving VISTA zero-points and in
comparing deeps to intermediates.

ETWS reported installing the 3rd party software on Nytro.


Non-survey Data Release:

None this time.
ACTION: MAR to supply list of non-surveys that could be released.


Astrogrid deployment, VO & Data Analysis services:

Nothing to report


Miscellaneous:

RGM noted that will be archiving the WFCAM raw data tapes from UKIRT.
-- 
Scanned by iCritical.